1.In a class of 30 students, 20 students like mathematics, 14 students like computer science, 6 students like neither. How many students like both mathematics and computer science?
Answer by ikleyn(52805)

.
1. In a class of 30 students, 20 students like mathematics, 14 students like computer science, 6 students like neither.
How many students like both mathematics and computer science?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

20 + 14 - x = 30-6,

34 - x = 24,

x = 34 - 24 = 10.

Answer.  10 students like both mathematics and computer science.

For details see the lesson
    - Counting elements in sub-sets of a given finite set
in this site.